### v1

# DeepTopoNet: A Framework for Subglacial Topography Estimation on the Greenland Ice Sheets

Abstract.
Understanding Greenland’s subglacial topography is critical for projecting the future mass loss of the ice sheet and its contribution to global sea-level rise. However, the complex and sparse nature of observational data, particularly information about the bed topography under the ice sheet, significantly increases the uncertainty in model projections. Bed topography is traditionally measured by airborne ice-penetrating radar that measures the ice thickness directly underneath the aircraft, leaving data gap of tens of kilometers in between flight lines. This study introduces a deep learning framework, which we call as DeepTopoNet, that integrates radar-derived ice thickness observations and BedMachine Greenland data through a novel dynamic loss-balancing mechanism. Among all efforts to reconstruct bed topography, BedMachine has emerged as one of the most widely used datasets, combining mass conservation principles and ice thickness measurements to generate high-resolution bed elevation estimates. The proposed loss function adaptively adjusts the weighting between radar and BedMachine data, ensuring robustness in areas with limited radar coverage while leveraging the high spatial resolution of BedMachine predictions i.e. bed estimates. Our approach incorporates gradient-based and trend surface features to enhance model performance and utilizes a convolutional neural network (CNN) architecture (i.e., BedTopoCNN) designed for subgrid-scale predictions. By systematically testing on the Upernavik Isstrøm) region, the model achieves high accuracy (MAE: 12.49, RMSE: 19.38, and R2: 0.99), outperforming baseline methods in reconstructing subglacial terrain. This work demonstrates the potential of deep learning in bridging observational gaps, providing a scalable and efficient solution to inferring subglacial topography. The framework offers implications for ice sheet modeling, paving the way for improved predictions of ice sheet flow and their impact on sea level rise.
1. Introduction
The Greenland Ice Sheet is a major contributor to global sea-level rise, driven by dynamic processes such as ice flow and iceberg calving. Rising sea levels pose a critical threat to coastal communities worldwide, increasing the risk of flooding, shoreline erosion, and habitat loss (Team et al., 2023). Estimating these contributions requires a detailed understanding of subglacial topography, which influences ice flow and basal friction (Barnes et al., 2021; Åkesson et al., 2021). However, accurately inferring the bed topography remains a significant challenge due to limited observational coverage, especially in the ice sheet interior. Traditional numerical methods, such as mass conservation models and geostatistic techniques (Morlighem et al., 2011; Bamber et al., 2001, 2013), have been instrumental but often struggle with sparse data.
Deep learning (LeCun et al., 2015), with its ability to learn complex relationships from multimodal data, has shown promise in solving geophysical problems. In this study, we present a deep learning framework for inferring Greenland’s subglacial topography, integrating radar-derived observations and BedMachine (Morlighem et al., 2017, 2014) data. A key innovation in our approach is the introduction of a dynamic loss-balancing mechanism that adaptively adjusts the contribution of radar and BedMachine data to the training process. This ensures robust performance in regions with sparse radar data while leveraging the high-resolution predictive capabilities of BedMachine.
Our methodology involves a convolutional neural network (CNN) architecture (LeCun et al., 2015) designed to capture spatial patterns in surface and subglacial features. By incorporating gradient-based covariates and trend-surface features, the model enhances its capacity to learn subgrid-scale variations and reduces the uncertainty in bed topography prediction. Additionally, the dynamic loss function ensures that the model optimally balances data-driven predictions and physical constraints derived from multiple data sources.
In this paper, we evaluate our proposed framework using a subset of Greenland’s bed topography. Our results demonstrate the model’s ability to accurately reconstruct subglacial terrain, outperforming traditional interpolation methods and baseline machine learning approaches. This work highlights the potential of deep learning in addressing critical geophysical challenges and provides a scalable solution for improving ice sheet models. The insights gained from this study contribute to a deeper understanding of subglacial dynamics and their implications for sea-level rise projections.
To summarize, the contributions of this paper are threefold: (i) DeepTopoNet Framework: A novel deep learning framework integrating radar observations and BedMachine data, featuring a CNN architecture enhanced with gradient-based covariates and trend-surface features; (ii) Dynamic Loss-Balancing Mechanism: An adaptive mechanism ensuring robust performance in regions with sparse radar data by balancing contributions from multimodal data sources; and (iii) Comprehensive Evaluation: Extensive benchmarking across Greenland’s Upernavik and unseen regions, demonstrating superior performance over traditional interpolation and machine learning methods.
2. Related Work
Spatial Interpolation. Spatial interpolation (Li et al., 2000; Lam, 1983) is a crucial methodology for estimating values at unobserved locations based on spatial data and has wide applications in hydrology, meteorology, and geosciences. Traditional interpolation techniques include deterministic methods, such as Inverse Distance Weighting (IDW) (Li and Heap, 2008) and Triangulated Irregular Network (TIN) (Feng et al., 2024), which rely on predefined spatial correlation functions. Geostatistical methods (Wackernagel and Wackernagel, 2003a, b), like Kriging and its variants (e.g., Ordinary Kriging and Universal Kriging), use spatial statistics and variogram models to infer underlying spatial patterns (Bamber et al., 2013). However, these methods are often limited by their reliance on fixed assumptions about spatial correlations and Gaussian process assumptions, which may not hold for many real-world datasets. Moreover, they rely solely on bed topography observations, which are sparse, and do not utilize any other ancillary dataset which may help guide the interpolation (e.g., deep trough generally coincide with observed high ice flow speed).
Recent developments in machine learning have enabled more sophisticated approaches to spatial interpolation. For example, Kriging Convolutional Networks (KCN) (Appleby et al., 2020) integrate the strengths of graph neural networks (GNNs) (Kipf and Welling, 2017) and kriging to enhance spatial data modeling. KCN leverages GNN-based propagation mechanisms while maintaining kriging-like inductive capabilities, offering flexibility and improved accuracy over standard kriging techniques. Similarly, the Graph for Spatial Interpolation (GSI) model adapts GNN architectures to dynamically learn spatial correlations while addressing limitations of fixed adjacency matrices, demonstrating effectiveness for rainfall spatial interpolation (Li et al., 2023). These approaches underscore the growing role of machine learning in addressing challenges such as irregular spatial patterns and dynamic spatial dependencies.
Subglacial Bed Topography Prediction. Inferring the subglacial bed topography has been a critical focus in glaciology due to its implications for ice sheet mass balance and their contributions to global sea-level rise. Recent advancements have integrated machine learning, physics-based models, and hybrid approaches to tackle this complex problem. Cheng et al. (2024) (Cheng et al., 2024) applied Physics-Informed Neural Networks (PINNs) to model ice flow dynamics for Helheim Glacier in Greenland. The study demonstrated PINNs’ capability to infer basal friction coefficients and interpolate sparse ice thickness observations with high accuracy. The framework uniquely incorporates the conservation of momentum (i.e., the shallow-shelf/shelfy-stream approximation) (MacAyeal, 1989; Morland, 1987) as soft constraints in the loss function, improving predictions where data are sparse and enhancing the versatility of the model. Yi et al. (2023) (Yi et al., 2023) evaluated nine machine learning and statistical models, including XGBoost, Gaussian Process Regression, and hybrid models, to predict Greenland’s subglacial bed topography. The study highlighted the strong performance of XGBoost when paired with kriging-based interpolation, achieving the best RMSE and R2. Similarly, Leong and Horgan (2020) (Leong and Horgan, 2020) introduced DeepBedMap, a deep learning framework based on generative adversarial networks, to produce high-resolution bed elevation maps of Antarctica. Their approach combined multi-resolution spatial data, such as surface elevation and ice velocity, to improve the representation of subglacial terrain roughness. These studies underscore the importance of combining domain knowledge with advanced computational tools to address challenges in regions with sparse observations. This paper builds on these foundations, proposing a robust framework that integrates multimodal data, deep learning, and loss balancing to further enhance predictive performance.
3. Preliminary
Accurately inferring Greenland’s subglacial topography is essential for understanding ice dynamics and predicting sea-level rise. However, this task is challenging due to sparse radar observations, limited direct measurements, and the inherent complexity of the ice-bed interactions. The goal of this paper is to develop a predictive framework that leverages surface observations and physical constraints to estimate bed topography, integrating multiple data sources. Let be the bed elevation at coordinates . Given sparse radar measurements and the modeled topography data from BedMachine , the objective is to predict with high spatial resolution across the domain . This can be formulated as:
where is surface elevation, is surface ice velocity components, is surface thickening rate over time, is surface mass balance (e.g., snow accumulation or surface melt), is gradients of all input features (e.g., surface elevation, surface velocity, surface thickening rate, and SMB), and is trend surface features generated from polynomial fitting of spatial coordinates.
4. Proposed Framework
The proposed framework, DeepTopoNet, integrates radar and BedMachine data to infer Greenland’s subglacial topography. Gradient and trend surface features enhance spatial variability capture, and the data is processed through BedTopoCNN, a convolutional neural network with residual blocks. A dynamic loss function balances contributions from radar and BedMachine data for robust predictions. The overall framework is shown in Figure 1.
4.1. Prior Physical Model (BedMachine)
BedMachine (Morlighem et al., 2017, 2020) is the most widely used subglacial topography map of Greenland. It employs a mass conservation approach in fast flowing regions, combining radar-derived ice thickness measurements, satellite velocity data, and surface mass balance reconstructions to produce a high-resolution bed topography map. The idea of mass conservation is to infer the bed elevation by ensuring conservation of mass: , where is the ice velocity, is surface elevation, is bed elevation, and is the apparent mass balance that combines observed thickness change and surface mass balance from a regional climate model. BedMachine ensures physically realistic reconstructions in regions of fast-flowing ice, but remains highly uncertain in slower moving regions where it is based on a simple interpolation scheme that connects existing radar data. The output is a seamless topographic map with horizontal resolution up to 150 m. In this study, BedMachine data serves as a reference for the target variable, subglacial bed topography (). The proposed DeepTopoNet framework goes beyond the interpolation based approach and the reliance on mass conservation.
4.2. Input Features and Covariates
To enhance predictive accuracy, the model incorporates both raw data and derived features: (i) Gradient Features (Jacobs, 2005). The gradient of raw input features X is computed as partial derivatives along - and -directions: , where , . These gradients provide additional spatial information on the variability of each feature and are concatenated to the input feature set. (ii) Trend Surface Features (Agterberg, 2023). For each input feature X, trend surface features are derived by fitting a polynomial of degree over the spatial coordinates: , where are coefficients estimated for X. These trend surfaces capture global spatial patterns and improve the model’s ability to generalize in regions with sparse data.
The integration of gradient features and trend surface features into DeepTopoNet represents a novel approach for subglacial topography estimation. Gradient features enhance local spatial sensitivity by capturing fine-scale variations in , , , and , while trend surface features provide a global spatial context by modeling large-scale variations using polynomial fitting. While these techniques have been used individually in geostatistics (Briggs, 1974; Hutchinson, 1995; Wang et al., 2023; Merriam and Lippert, 1966) and image processing (Riba et al., 2020; Canny, 1986; Sobel, 1970; Soria et al., 2023), their combined use in a deep learning framework for sparse spatial predictions is unprecedented. Our hybrid approach improves the model’s ability to generalize in regions with sparse observations, bridging local feature variability with global spatial trends to enhance topographic reconstruction accuracy.
4.3. Model Architecture
The proposed model architecture, BedTopoCNN, shown in Figure 2 is a deep convolutional neural network specifically designed for subglacial topography prediction. Building on residual learning principles introduced by He et al. (2016) (He et al., 2016), the architecture employs a series of Residual Blocks (RBs) to capture spatial features from multimodal input data effectively while addressing challenges such as vanishing gradients during training. Each RB consists of two convolutional layers with a kernel size of , padding, and batch normalization to ensure feature stability. Non-linear transformations are applied using ReLU activation, while a skip connection (adapted from (He et al., 2016)) allows the input to bypass the convolutional layers, preserving features and mitigating gradient vanishing issues. Additionally, Dropout regularization is incorporated within the blocks to prevent overfitting, which is particularly important in regions with sparse radar data (Cai et al., 2025; Leong and Horgan, 2020).
The overall architecture begins with an Input Layer that processes the preprocessed input features. This is followed by five RBs, where the number of filters progressively increases from 32 to 256 to extract hierarchical spatial features. Finally, the network concludes with a convolutional layer that reduces the output to a single channel, corresponding to the predicted bed topography. The combination of RBs and the hierarchical design allows the network to effectively model complex relationships between input features and the target subglacial topography.
4.4. Loss Function
We propose a dynamic loss function to guide the training process by balancing the contributions of radar data and BedMachine data . While dynamic loss balancing has been explored in multi-task learning (Kendall et al., 2018) and semi-supervised settings (Wang et al., 2021), our approach uniquely adapts to geospatial data sparsity patterns rather than task uncertainty or pseudo-label confidence. Furthermore, in contrast to PINNs, where loss weights are often manually tuned or fixed (Raissi et al., 2019), our method dynamically adjusts the weighting terms and based on the local density of radar observations in each batch. The total loss is expressed as , where:
is the radar loss, penalizing discrepancies in radar-covered regions ,
is the BedMachine loss, evaluating prediction accuracy in non-radar regions , and ,
are dynamic weights that adjust during training, ensuring balanced optimization with as a small smoothing factor to avoid division by zero.
The objective is to train a a deep convolutional neural network model with parameters to minimize the total loss , ensuring robust predictions across radar-sparse and radar-dense regions: .
5. Experiments
In this section, we evaluate the accuracy of DeepTopoNet and demonstrate its predictive capability across the entire grid map.
5.1. Experimental Setup
5.1.1. Dataset
The dataset used in this study is obtained from the Upernavik Isstrøm region in the Central West of Greenland, encompassing four distinct sub-regions. Each sub-region covers a spatial extent of grid cells, representing an area with diverse ice dynamics and varying data density. The number of radar data points differs across the sub-regions, reflecting the uneven distribution of observations. The dataset includes several observed variables critical for subglacial topography inference:
: Surface elevation (Howat et al., 2014), representing the height of the ice sheet surface above a reference level.
: Surface ice velocity components (Joughin et al., 2010) in the - and -directions, describing ice flow dynamics.
: Surface thickening or thinning rate (Millan et al., 2022), indicating changes in ice sheet thickness over time.
These variables provide essential input features for the model, capturing spatial and temporal variations in surface and subsurface ice dynamics. The dataset’s diverse sub-regions allow the DeepTopoNet to generalize across areas with different radar data availability and ice sheet characteristics, enhancing the robustness of subglacial topography predictions.
The spatial relationship between radar observations and BedMachine topography data is critical for understanding the integration of data in subglacial bed topography modeling. A detailed visualization of the BedMachine data with overlaid radar observation points for each sub-region in the Upernavik Isstrøm region is provided in Figure 3. This figure highlights the spatial distribution of radar data relative to the underlying bed topography, showcasing the coverage of radar tracks and emphasizing the importance of filling data gaps in regions with sparse radar measurements.
| Sub-regions | Min_bed_height | Max_bed_height | #Radar data points |
| I | -478.7407 | 778.9677 | 316,559 |
| II | -144.0005 | 646.0179 | 238,899 |
| III | 115.8914 | 762.7296 | 19,509 |
| IV | 106.4424 | 428.6476 | 57,340 |
5.1.2. Evaluation Metrics
The model’s performance is assessed during training and validation, as well as through a comprehensive evaluation on the full prediction grid. For the evaluation of full grid predictions, several standard metrics are employed to comprehensively measure the model’s accuracy and structural fidelity such as MAE, RMSE, Structural Similarity Index Measure (SSIM) (Wang et al., 2004), and Peak Signal-to-Noise Ratio (PSNR) (Hore and Ziou, 2010). These metrics, commonly used in spatial modeling and image reconstruction tasks, provide a holistic evaluation of the model’s accuracy, structural preservation, and robustness for full grid predictions.
Moreover, a domain-specific metric, the Terrain Ruggedness Index (TRI) (Reily Shawn et al., 1999), is used as a quantitative measure to evaluate local variations in terrain elevation by capturing the differences between a cell and its neighboring elevation values. TRI provides insight into how well a model preserves the spatial variability and ruggedness of the terrain, which is critical for applications such as glaciological modeling and subglacial flow analysis. By comparing TRI values between the predicted topography and the reference BedMachine data, we assess the model’s ability to reproduce terrain features accurately. For a given digital elevation model (DEM), TRI is calculated for each grid cell by measuring the elevation differences between the cell and its neighboring cells. The value represents the root mean square of these differences, capturing the local terrain variability. To evaluate the model predictions, we leverage relative difference (%), which is the relative difference that quantifies how closely the ruggedness of the predicted topography aligns with the reference BedMachine data:
where a lower relative difference indicates better alignment between the predicted and reference ruggedness.
Algorithm 1 describes the calculation of TRI for a grid-based Digital Elevation Model (DEM). For each cell in the elevation grid, the algorithm calculates the squared differences in elevation between the cell and its neighboring cells (defined by offsets ). These differences are accumulated across all grid cells and normalized by the total number of neighboring pairs. The TRI is then computed as the square root of the mean squared elevation differences, providing a measure of terrain variability. This approach captures fine-scale topographic details, making it useful for characterizing subglacial landscapes.
| Method | Sub-region I | Sub-region II | Sub-region III | Sub-region IV | ||||||||||||||||
| MAE | RMSE | R | SSIM | PSNR | MAE | RMSE | R | SSIM | PSNR | MAE | RMSE | R | SSIM | PSNR | MAE | RMSE | R | SSIM | PSNR | |
| IDW (Johnston et al., 2001) | 63.15 | 101.05 | 0.65 | 0.78 | 21.90 | 29.90 | 41.84 | 0.82 | 0.87 | 25.52 | 52.16 | 74.92 | 0.58 | 0.91 | 18.72 | 22.99 | 34.31 | 0.67 | 0.94 | 19.46 |
| RBF (Heryudono and Driscoll, 2010) | 65.95 | 99.93 | 0.66 | 0.77 | 22.00 | 35.28 | 53.38 | 0.70 | 0.85 | 23.40 | 57.48 | 76.52 | 0.56 | 0.90 | 18.54 | 20.56 | 26.72 | 0.80 | 0.94 | 21.63 |
| RF (Sekulić et al., 2020) | 68.00 | 99.87 | 0.66 | 0.69 | 22.00 | 79.90 | 97.13 | 0.01 | 0.59 | 18.21 | 61.90 | 93.35 | 0.34 | 0.84 | 16.81 | 24.46 | 34.13 | 0.67 | 0.84 | 19.50 |
| MLP-PE (Li et al., 2021) | 113.17 | 195.27 | -0.30 | 0.48 | 16.18 | 112.51 | 170.17 | -2.03 | 0.28 | 13.34 | 145.43 | 238.89 | -3.30 | 0.41 | 8.65 | 88.49 | 151.62 | -5.46 | 0.19 | 6.55 |
| VAE+XGBoost (Yi et al., 2023) | 78.12 | 108.54 | 0.60 | 0.75 | 21.28 | 42.52 | 59.79 | 0.63 | 0.75 | 22.42 | 57.91 | 85.77 | 0.45 | 0.85 | 17.55 | 22.95 | 30.88 | 0.73 | 0.83 | 20.37 |
| U-Net (Ronneberger et al., 2015) | 56.43 | 75.59 | 0.81 | 0.82 | 24.42 | 46.81 | 68.57 | 0.51 | 0.72 | 21.23 | 69.83 | 106.39 | 0.15 | 0.64 | 15.68 | 14.17 | 20.91 | 0.88 | 0.88 | 23.76 |
| U-Net++ (Zhou et al., 2018) | 88.36 | 114.47 | 0.55 | 0.78 | 20.82 | 41.83 | 65.24 | 0.55 | 0.76 | 21.66 | 61.42 | 94.61 | 0.33 | 0.68 | 16.70 | 9.62 | 14.26 | 0.94 | 0.88 | 27.08 |
| U-Net3+ (Huang et al., 2020) | 51.33 | 70.92 | 0.83 | 0.77 | 24.98 | 33.78 | 47.82 | 0.76 | 0.80 | 24.36 | 62.94 | 97.77 | 0.28 | 0.72 | 16.41 | 24.95 | 34.99 | 0.66 | 0.88 | 19.28 |
| Att. U-Net (Oktay et al., 2018) | 21.88 | 33.33 | 0.96 | 0.90 | 31.53 | 21.35 | 33.21 | 0.89 | 0.87 | 27.53 | 20.65 | 31.54 | 0.93 | 0.90 | 26.24 | 7.33 | 11.75 | 0.96 | 0.94 | 28.76 |
| DeepTopoNet | 12.49 | 19.38 | 0.99 | 0.94 | 36.25 | 18.26 | 28.90 | 0.91 | 0.88 | 28.73 | 21.00 | 32.69 | 0.92 | 0.91 | 25.93 | 4.74 | 7.65 | 0.98 | 0.96 | 32.49 |
| Method | Sub-region I | Sub-region II | Sub-region III | Sub-region IV | ||||||||||||||||
| MAE | RMSE | R | SSIM | PSNR | MAE | RMSE | R | SSIM | PSNR | MAE | RMSE | R | SSIM | PSNR | MAE | RMSE | R | SSIM | PSNR | |
| w/o stride,, & t | 42.84 | 62.66 | 0.87 | 0.80 | 26.05 | 28.91 | 41.00 | 0.82 | 0.85 | 25.70 | 22.44 | 34.52 | 0.91 | 0.91 | 25.46 | 87.99 | 457.06 | -57.74 | 0.88 | -3.04 |
| w/o stride, w/ & t | 23.27 | 36.26 | 0.96 | 0.88 | 30.80 | 18.42 | 28.78 | 0.91 | 0.89 | 28.77 | 17.37 | 27.55 | 0.94 | 0.90 | 27.41 | 4.80 | 7.53 | 0.98 | 0.96 | 32.63 |
| w/o | 39.21 | 67.38 | 0.85 | 0.79 | 25.42 | 34.44 | 54.09 | 0.69 | 0.73 | 23.29 | 65.13 | 104.42 | 0.18 | 0.66 | 15.84 | 92.42 | 106.00 | -2.16 | 0.61 | 9.66 |
| DeepTopoNet | 12.49 | 19.38 | 0.99 | 0.94 | 36.25 | 18.26 | 28.90 | 0.91 | 0.88 | 28.73 | 21.00 | 32.69 | 0.92 | 0.91 | 25.93 | 4.74 | 7.65 | 0.98 | 0.96 | 32.49 |
5.1.3. Baselines
We compare the performance of the proposed DeepTopoNet against three categories of bed topography prediction baselines, which include traditional interpolation methods: IDW (Johnston et al., 2001) and RBF (Heryudono and Driscoll, 2010); pixel-based machine learning methods: random forest (RF) (Sekulić et al., 2020), MLP with positional encoding (MLP-PE) (Li et al., 2021), and VAE with XGBoost (Yi et al., 2023); and patch-based deep learning models: U-Net (Ronneberger et al., 2015), U-Net++ (Zhou et al., 2018), U-Net3+ (Huang et al., 2020), and Attention U-Net (Oktay et al., 2018). Below is a brief explanation of each method:
IDW (Johnston et al., 2001): A deterministic interpolation technique that assigns weights to nearby points inversely proportional to their distance from the target point. This method emphasizes closer points for prediction, assuming spatially smooth variations.
RBF (Heryudono and Driscoll, 2010): An interpolation method that uses radial basis functions to create a smooth surface fitting the data. RBF is known for its flexibility in capturing localized spatial variations.
RF (Sekulić et al., 2020): This baseline utilizes an RF regressor with additional spatial enhancements to predict bed topography.
MLP with positional encoding (MLP-PE) (Li et al., 2021): This baseline employs a Multilayer Perceptron (MLP) model enhanced with sinusoidal positional encodings to incorporate spatial context into the predictions.
VAE with XGBoost (Yi et al., 2023): This baseline leverages a Variational Autoencoder (VAE) to encode spatial patterns into a latent representation, followed by an XGBoost regressor to predict subglacial topography with enhanced feature extraction.
U-Net (Ronneberger et al., 2015): A widely used convolutional neural network for segmentation tasks, featuring an encoder-decoder structure with skip connections to preserve spatial details.
U-Net++ (Zhou et al., 2018): An extension of U-Net that introduces dense skip connections to improve feature aggregation and capture finer details.
U-Net3+ (Huang et al., 2020): A further enhancement of U-Net with nested and multi-scale skip connections, designed to leverage information across different feature resolutions.
Attention U-Net (Oktay et al., 2018): An advanced U-Net variant that incorporates attention mechanisms to focus on critical regions of the input, enhancing the model’s ability to capture important spatial features.
5.1.4. Experimental Settings
During training, the dataset is split into 80% for training and 20% for validation, ensuring an unbiased evaluation of the model’s generalization capabilities. The training data is processed in overlapping patches of size with a stride of 8, increasing the effective dataset size and capturing both local and contextual spatial features. The model is trained with a batch size of 16 using the Adam optimizer for 20,000 epochs. A cyclic learning rate scheduler dynamically adjusts the learning rate during training. The loss function dynamically balances radar and BedMachine data contributions. Early stopping is employed with a patience threshold of 5,000 epochs. The best model is saved when the validation loss improves, ensuring optimal performance and preventing overfitting. We present hyperparameter settings for each method as follows.
IDW. Number of Nearest Neighbors (): 4000, power parameter (): 2, distance threshold = 1010.
RBF. The function is used to capture smooth variations in the data, smoothness parameter (): 2, small regularization term: .
Random forest. The hyperparameter settings were optimized using Optuna (Akiba et al., 2019), a framework for automated hyperparameter optimization. These hyperparameters were tuned as follows. : 1001000, : 210, : 110, : . The objective function optimized the RMSE on the test set. The optimization process was conducted using 100 trials with the TPE (Tree-structured Parzen Estimator) sampler for efficient exploration of the search space.
MLP with Positional Encoding (MLP-PE). Sinusoidal positional encodings are added to the input data to incorporate spatial context, with : 15. The input features are augmented using gradient covariates, polynomial trend surface, and nearest neighbor values using = 3000. MLP architecture consist of two fully connected layers with 64 hidden units each and ReLU is applied after each hidden layer. Adam optimizer with a learning rate of is used, where MSE loss evaluated the difference between predicted and true values during training. We use batch size of 64, and the model was trained for 20,000 epochs with an early stopping patience: 5,000 epochs.
VAE with XGBoost. The VAE encoder consists of three fully connected layers (32, 64, and 16 neurons) with a latent dimension of 3, followed by a decoder with a mirrored structure. ReLU activation is applied after each hidden layer. The model is trained using MSE reconstruction loss and KL divergence, optimized with Adam () learning rate) for 100 epochs with early stopping (patience: 50 epochs). Latent features extracted from the trained VAE are used as input for XGBoost, which undergoes hyperparameter tuning with Optuna over 100 trials, optimizing tree depth (312), learning rate (0.010.3), and regularization terms. The final model is trained using histogram-based optimization (tree_method=’hist’) with GPU acceleration.
U-Net. The architecture of U-Net is as follows. Encoder: 4 convolutional blocks with 64, 128, 256, and 512 filters, each block consisting of two convolutions, batch normalization, and ReLU, followed by max pooling. Bottleneck: A 1024-filter convolutional block. Decoder: Four up-convolutional blocks for upsampling, each followed by concatenation with corresponding encoder features and convolutional layers. The number of filters reduces progressively from 512 to 64. Output layer: A convolutional layer maps the 64-filter feature map to a single output channel. Training settings was used as follows: Adam optimizer with a learning rate of and weight decay of , Cyclic learning rate strategy with = 10-5 and = 10-3, batch size: 16, number of epoch: 20K with early stopping triggered after 5K epochs.
U-Net++. The architecture of U-Net++ is as follows. Encoder Path: Four levels of convolutional layers with increasing channels (64, 128, 256, 512), followed by max-pooling layers for down-sampling. Decoder Path: Decoding layers with up-sampling (bilinear interpolation) and skip connections from corresponding encoder levels, progressively reducing the spatial dimension and channel count. Final layer: A convolution reduces the output to a single channel representing predicted topography. Training settings were the same as the original U-Net.
U-Net3+. The architecture of U-Net3+ is as follows. Encoder Path: Four levels of convolutional layers, where the number of channels increases progressively (64, 128, 256, 512). Each level consists of a convolutional block with a kernel, batch normalization, and ReLU activation, followed by max-pooling layers for down-sampling. Decoder Path with Full-Scale Skip Connections: The decoder utilizes full-scale skip connections, where features from all previous encoder levels are concatenated with the current decoder level after up-sampling (using bilinear interpolation). This structure allows multi-scale features to be effectively combined at each decoding stage. The number of channels is progressively reduced (512 to 256, 256 to 128, and 128 to 64) using convolutional layers with batch normalization and ReLU activation. Final Layer: A convolutional layer reduces the output to a single channel representing the predicted bed topography. Training settings were the same as the original U-Net.
Attention U-Net. The architecture of Attention U-Net is as follows. Encoder Path: Four levels of convolutional layers with increasing channels (64, 128, 256, 512), where each level consists of two convolutional layers followed by ReLU activation. Down-sampling is performed using max-pooling layers, progressively reducing the spatial dimensions while increasing feature depth. Decoder Path: Symmetrical to the encoder, the decoder performs up-sampling using transposed convolution layers. Each up-sampled feature map is concatenated with the corresponding encoder feature map through skip connections, enhanced by Attention Blocks. The attention mechanism selectively focuses on relevant spatial regions by weighting the encoder features based on their interaction with the decoder features. Attention Mechanism: The attention blocks compute weights using gating signals () and skip-connected encoder features (), applying a learned attention mask to the encoder features. Final Layer: A convolutional layer reduces the output to a single channel representing the predicted topography. The training settings are the same as the original U-Net.
5.2. Results
In this section, we present the evaluation results of the proposed framework, including a quantitative comparison across multiple performance metrics, an ablation study to assess the contributions of key components, a qualitative comparison for visual analysis of the predictions, and spatial generalization analysis. All experiments were conducted on a PC with Linux OS, Intel Xeon 2.4GHz processor, 256GB RAM, and GPU RTX A4500.
5.3. Quantitative Comparison
The proposed framework, DeepTopoNet, demonstrates superior performance across all sub-regions in the Upernavik Isstrøm region. Comprehensive evaluation metrics, including MAE, RMSE, R2, SSIM, and PSNR, consistently show that DeepTopoNet achieves the lowest errors and the highest similarity indices across all sub-regions. This highlights its robustness in accurately predicting bed topography, even in regions with varying radar data density. Table 2 provides a detailed quantitative comparison of all methods across the four sub-regions.
Compared to traditional interpolation methods, such as IDW (Johnston et al., 2001) and RBF (Heryudono and Driscoll, 2010), DeepTopoNet delivers significantly improved performance. For instance, while IDW achieves moderate results in some sub-regions (e.g., SSIM of 0.87 in Sub-region II), its RMSE values remain substantially higher (e.g., 101.05 in Sub-region I). Similarly, RBF fails to capture finer spatial details, resulting in higher errors across all metrics. In contrast, DeepTopoNet achieves considerably lower RMSE and higher SSIM, demonstrating its ability to model spatial variability more effectively than interpolation methods.
Against pixel-based machine learning methods, including RF (Sekulić et al., 2020), MLP-PE (Li et al., 2021) and VAE+XGBoost (Yi et al., 2023), DeepTopoNet exhibits a significant advantage. RF struggles to capture complex spatial relationships, as reflected in its relatively high RMSE values (e.g., 99.87 in Sub-region I) and low R2 scores across sub-regions. MLP-PE, despite incorporating positional encoding, performs poorly with high errors and negative R2 in some sub-regions (e.g., Sub-region III), indicating limited generalization. In contrast, DeepTopoNet effectively leverages spatial features and multimodal data, achieving an RMSE as low as 19.38 and an SSIM of 0.94 in Sub-region I. VAE+XGBoost performs better than MLP-PE and RF but still falls short of DeepTopoNet in all sub-regions. While VAE+XGBoost benefits from variational autoencoder-based feature extraction and gradient-boosted decision trees for regression, its RMSE remains significantly higher than DeepTopoNet across all sub-regions (e.g., 108.54 in Sub-region I vs. 19.38 for DeepTopoNet).
When compared to patch-based deep learning models, including U-Net (Ronneberger et al., 2015), U-Net++ (Zhou et al., 2018), U-Net3+ (Huang et al., 2020), and Attention U-Net (Oktay et al., 2018), DeepTopoNet shows a clear improvement. Among these baselines, Attention U-Net performs the best, achieving competitive results with an RMSE of 33.33 and SSIM of 0.90 in Sub-region I. However, DeepTopoNet consistently surpasses Attention U-Net, achieving a much lower RMSE of 19.38 and a higher SSIM of 0.94 in the same sub-region. This indicates that DeepTopoNet’s combination of enriched input features, dynamic loss balancing, and robust architecture allows it to better capture spatial relationships and generalize effectively across diverse sub-regions. In summary, DeepTopoNet’s consistent outperformance across all categories of baselines demonstrates its superior capability for accurately reconstructing subglacial topography, providing a robust solution for regions with complex and sparse observational data.
Table 4 presents the relative difference (%) in TRI for DeepTopoNet and baseline methods, assessing their ability to capture terrain variability. Traditional interpolation methods, such as IDW and RBF, perform well in less rugged areas like Sub-region IV but struggle in complex terrains like Sub-region II, where RBF exhibits a high TRI deviation of 57.02%. Similarly, pixel-based models, including Random Forest and MLP-PE, show significant TRI deviations, with MLP-PE exceeding 883.57% in Sub-region II and 1223.73% in Sub-region IV, indicating their inability to effectively model spatial dependencies and terrain complexity.
| Method | Sub-region I | Sub-region II | Sub-region III | Sub-region IV |
| IDW (Johnston et al., 2001) | 25.71 | 32.73 | 8.13 | 3.55 |
| RBF (Heryudono and Driscoll, 2010) | 32.20 | 57.02 | 15.77 | 1.98 |
| RF (Sekulić et al., 2020) | 85.33 | 214.49 | 35.03 | 26.07 |
| MLP-PE (Li et al., 2021) | 451.44 | 883.57 | 583.15 | 1223.73 |
| VAE+XGBoost (Yi et al., 2023) | 23.05 | 97.50 | 24.22 | 36.12 |
| U-Net (Ronneberger et al., 2015) | 19.26 | 97.59 | 230.51 | 2.50 |
| U-Net++ (Zhou et al., 2018) | 9.42 | 59.34 | 150.98 | 6.74 |
| U-Net3+ (Huang et al., 2020) | 56.56 | 72.90 | 143.71 | 5.37 |
| Att. U-Net (Oktay et al., 2018) | 18.05 | 25.11 | 18.06 | 10.68 |
| DeepTopoNet | 6.58 | 30.32 | 8.01 | 1.20 |
Among patch-based deep learning models, U-Net variants demonstrate varying performance, with U-Net++ performing well in Sub-regions I and IV but struggling in Sub-region III (150.98%), suggesting difficulty in adapting to different terrain structures. Attention U-Net achieves strong results in Sub-region II (25.11%) but is inconsistent across other regions. DeepTopoNet consistently achieves the lowest TRI differences in Sub-regions I, III, and IV (6.58%, 8.01%, and 1.20%, respectively), demonstrating its robustness in preserving terrain ruggedness. While its performance in Sub-region II (30.32%) leaves room for improvement, it remains competitive with the best-performing baselines. These results confirm that DeepTopoNet effectively generalizes across diverse landscapes while maintaining superior topographic accuracy.
5.4. Ablation Study
To further assess the contribution of each component in DeepTopoNet, we conducted an ablation study by systematically removing components to evaluate their impact on the model’s performance. The study aimed to quantify how the exclusion of these components affects predictive accuracy. First, to examine the importance of using patches with stride, gradient features, and trend surface features, we removed all three components. In this case, the model is trained using only the original 5 raw features without applying stride during patch extraction. Second, to isolate the effect of stride, we trained the model on patches without stride but retained the gradient and trend surface features. The results of the ablation experiments are summarized in Table 3.
w/o stride, gradient (), and trend surface (t) features. Removing these components leads to a substantial decline in performance. For instance, in Sub-region IV, RMSE increases drastically to 457.06, and R2 becomes negative (e.g., -57.74), indicating a failure to capture meaningful spatial patterns. In other sub-regions, metrics such as SSIM and PSNR are lower compared to DeepTopoNet, emphasizing the importance of these features for accurate predictions.
w/o Stride, w/ gradient (), and trend surface (t) features. Retaining gradient and trend surface features while removing stride significantly improves performance compared to the first scenario. For example, in Sub-region I, RMSE reduces to 36.26, and R2 increases to 0.96, demonstrating that gradient and trend surface features are vital for capturing spatial variability. However, the performance still lags behind DeepTopoNet in certain metrics, such as RMSE in Sub-region III (27.55 vs. 32.69 for DeepTopoNet).
w/o BedMachine Loss (). Removing the BedMachine Loss () significantly degrades the model’s performance, highlighting its critical role in training DeepTopoNet. Across all sub-regions, the absence of leads to substantial increases in RMSE, particularly in Sub-region IV, where RMSE jumps to 106.00 and R2 drops to -2.16, indicating a severe loss of predictive accuracy. Additionally, SSIM and PSNR values decline across all sub-regions, reflecting a reduction in structural similarity and topographic detail retention. The poor performance suggests that plays a crucial role in constraining the predictions, ensuring that the model maintains coherence with the global topographic trends derived from BedMachine data.
DeepTopoNet. The full DeepTopoNet framework achieves the best results across most metrics and sub-regions. For example, in Sub-region I, RMSE is the lowest at 19.38, and SSIM reaches 0.94, indicating excellent structural similarity between predictions and ground truth. This highlights the effectiveness of combining stride, , t, and for robust and accurate bed topography predictions. The ablation study shows that each component contributes significantly to the performance of DeepTopoNet. While and t play a key role in improving accuracy, the inclusion of stride further enhances the model’s ability to generalize and capture spatial context, leading to superior performance across all sub-regions. Additionally, the proves to be a crucial component of the model. This confirms that is essential for balancing high-resolution radar-derived observations with global-scale topographic trends, ensuring stability and accuracy in bed topography predictions.
| Prediction/BedMachine (Morlighem et al., 2017, 2020)/Difference |
| IDW (Johnston et al., 2001) |
| RBF (Heryudono and Driscoll, 2010) |
| RF (Sekulić et al., 2020) |
| MLP-PE (Li et al., 2021) |
| VAE+XGBoost (Yi et al., 2023) |
| U-Net (Ronneberger et al., 2015) |
| U-Net++ (Zhou et al., 2018) |
| U-Net3+ (Huang et al., 2020) |
| Att. U-Net (Oktay et al., 2018) |
| DeepTopoNet |
5.5. Qualitative Comparison
The qualitative comparison in Table 5 highlights the performance of various methods in predicting the full grid topography for Sub-region I. Each row presents the predicted topography, the reference BedMachine map, and the difference map, offering a visual assessment of spatial accuracy and consistency with ground truth.
Interpolation Methods (IDW and RBF). Both IDW and RBF provide smooth predictions but fail to capture finer spatial details, as evident from the difference maps. IDW tends to oversmooth areas with high variability, leading to notable discrepancies in regions of steep topography. Similarly, RBF produces smoother outputs but struggles with localized variations, particularly in areas influenced by sharp transitions in the reference map.
Pixel-Based Methods (RF, MLP-PE and VAE+XGB). Random forest (RF) shows improvement in capturing localized features compared to interpolation methods, yet the difference map reveals persistent inaccuracies in spatial patterns. On the other hand, MLP with positional encoding (MLP-PE) fails to generalize effectively, producing artifacts and noticeable errors in the difference map, reflecting its limited ability to model complex spatial relationships. VAE+XGBoost, which leverages a VAE to extract latent spatial representations before using XGBoost for prediction, exhibits moderate performance in reconstructing bed topography. While the VAE efficiently encodes high-dimensional spatial patterns, the combination with XGBoost still struggles in areas with sharp terrain transitions, leading to noticeable deviations in the difference map.
Patch-Based Methods (U-Net Variants). Among the U-Net family, standard U-Net provides better results compared to U-Net++ and U-Net3+, as evident from the more refined predictions and reduced discrepancies in the difference maps. Attention U-Net further improves spatial accuracy by focusing on critical regions, yielding smaller differences in areas of complex topography. However, all these methods exhibit slight smoothing in regions with abrupt elevation changes.
DeepTopoNet. The proposed DeepTopoNet demonstrates the best qualitative performance, with predictions closely matching the BedMachine reference. The difference map for DeepTopoNet shows minimal discrepancies, especially in areas with high variability and sharp transitions, highlighting its superior capability in capturing complex spatial patterns. This can be attributed to its robust architecture, dynamic loss balancing, and the integration of gradient and trend surface features, which enable precise predictions. Overall, DeepTopoNet outperforms all baseline methods in terms of spatial accuracy and alignment with the BedMachine reference. Its ability to generalize effectively across the entire grid and resolve fine spatial details underscores its robustness and reliability for subglacial topography prediction.
5.6. Spatial Generalization Analysis
To assess spatial generalization, we implement a grid-based validation strategy by horizontally slicing the input domain into 30 equal-width bands (Beery et al., 2018; Meyer et al., 2019). Odd-numbered slices (1, 3, 5, …) are used for training and even-numbered slices (2, 4, 6, …) for testing. This design mimics realistic extrapolation to unobserved transects, reflecting operational deployment where radar coverage is spatially sparse.
| Model | MAE | RMSE | R | SSIM | PSNR | TRI |
| U-Net | 65.92 | 90.52 | 0.72 | 0.79 | 22.86 | 15.24 |
| U-Net++ | 107.97 | 142.74 | 0.31 | 0.69 | 18.90 | 17.78 |
| U-Net3+ | 66.68 | 88.98 | 0.73 | 0.75 | 23.01 | 21.89 |
| Att. U-Net | 23.69 | 35.47 | 0.96 | 0.89 | 30.99 | 16.67 |
| DeepTopoNet | 15.90 | 25.39 | 0.98 | 0.91 | 33.90 | 17.94 |
We evaluate DeepTopoNet and representative deep learning baselines (e.g., U-Net, U-Net++, U-Net3+, and Attention U-Net) under this spatial holdout scheme. As shown in Table 6, DeepTopoNet achieves the best performance across all major evaluation metrics in Sub-region I. It produces the lowest MAE (15.90) and RMSE (25.39), while maintaining the highest (0.98), SSIM (0.91), and PSNR (33.90). Notably, even though U-Net and U-Net3+ perform moderately well, they fall short in both structural similarity and magnitude-based errors. While Attention U-Net achieves competitive performance, DeepTopoNet still outperforms it in all categories except TRI, where U-Net slightly edges out others in preserving localized ruggedness.
These results confirm that DeepTopoNet not only excels under standard training-validation splits but also maintains robust predictive performance when generalized to spatially unseen regions. This strengthens its applicability to real-world scenarios where radar data is geographically sparse and extrapolation is critical.
6. Conclusion
This paper introduces DeepTopoNet, a novel deep learning framework for predicting subglacial topography in regions with sparse radar data. By integrating multimodal features (e.g., gradients, trend surfaces) and employing a dynamic loss balancing mechanism, DeepTopoNet achieves superior accuracy in reconstructing Greenland’s bed topography across diverse sub-regions. Evaluations show it outperforms traditional interpolation, pixel-based machine learning, and state-of-the-art patch-based deep learning methods. Ablation studies confirm the importance of key components, such as stride, gradient features, and BedMachine loss, in enhancing model performance. DeepTopoNet excels in capturing fine spatial details and maintaining accuracy in varying data densities, demonstrating robustness and scalability. This framework not only improves subglacial topography modeling but also offers a scalable solution for other regions, bridging observational gaps and advancing insights into ice dynamics and sea-level rise. Furthermore, to support real-world deployment, we evaluate its spatial robustness via a grid-based slicing protocol that simulates extrapolation to unmeasured regions. In our future work, we aim to enhance the generalizability of DeepTopNet to varying subglacial topography by incorporating graph neural networks (GNNs) or knowledge-informed priors, which are better suited for capturing spatial dependencies and irregular topographic structures.
Acknowledgment
This research has been funded by the NSF HDR Institute for Harnessing Data and Model Revolution in the Polar Regions (iHARP), Award #2118285.
Code Availability
The source code for DeepTopoNet and all related experiments is available at: https://github.com/bayuat/DeepTopoNet

### v2

@@
 Abstract.
-Understanding Greenland’s subglacial topography is critical for projecting the future mass loss of the ice sheet and its contribution to global sea-level rise.
+Mapping Greenland’s subglacial topography is critical for projecting the future mass loss of the ice sheet and its contribution to global sea-level rise.
 However, the complex and sparse nature of observational data, particularly information about the bed topography under the ice sheet, significantly increases the uncertainty in model projections.
-Bed topography is traditionally measured by airborne ice-penetrating radar that measures the ice thickness directly underneath the aircraft, leaving data gap of tens of kilometers in between flight lines.
-This study introduces a deep learning framework, which we call as DeepTopoNet, that integrates radar-derived ice thickness observations and BedMachine Greenland data through a novel dynamic loss-balancing mechanism.
+Bed topography is traditionally measured by airborne ice-penetrating radars that measure the ice thickness directly underneath the aircraft, leaving data gaps of tens of kilometers in between flight lines.
+This study introduces a deep learning framework, DeepTopoNet, that integrates radar-derived ice thickness observations and BedMachine Greenland data through a novel dynamic loss-balancing mechanism.
 Among all efforts to reconstruct bed topography, BedMachine has emerged as one of the most widely used datasets, combining mass conservation principles and ice thickness measurements to generate high-resolution bed elevation estimates.
-The proposed loss function adaptively adjusts the weighting between radar and BedMachine data, ensuring robustness in areas with limited radar coverage while leveraging the high spatial resolution of BedMachine predictions i.e.
-bed estimates.
+The proposed loss function adaptively adjusts the weighting between radar and BedMachine’s bed, ensuring robustness in areas with limited radar coverage while leveraging the high spatial resolution of BedMachine’s bed estimates.
 Our approach incorporates gradient-based and trend surface features to enhance model performance and utilizes a convolutional neural network (CNN) architecture (i.e., BedTopoCNN) designed for subgrid-scale predictions.
-By systematically testing on the Upernavik Isstrøm) region, the model achieves high accuracy (MAE: 12.49, RMSE: 19.38, and R2: 0.99), outperforming baseline methods in reconstructing subglacial terrain.
+By systematically testing on the Upernavik Isstrøm) region in West Greenland, the model achieves high accuracy (MAE: 12.49 m, RMSE: 19.38 m, and R2: 0.99), outperforming baseline methods in reconstructing subglacial terrain.
 This work demonstrates the potential of deep learning in bridging observational gaps, providing a scalable and efficient solution to inferring subglacial topography.
-The framework offers implications for ice sheet modeling, paving the way for improved predictions of ice sheet flow and their impact on sea level rise.
+This framework paves the way for improved predictions of ice sheet flow and sea level rise.
 1. Introduction
-The Greenland Ice Sheet is a major contributor to global sea-level rise, driven by dynamic processes such as ice flow and iceberg calving.
-Rising sea levels pose a critical threat to coastal communities worldwide, increasing the risk of flooding, shoreline erosion, and habitat loss (Team et al., 2023).
-Estimating these contributions requires a detailed understanding of subglacial topography, which influences ice flow and basal friction (Barnes et al., 2021; Åkesson et al., 2021).
+The Greenland Ice Sheet is a major contributor to global sea-level rise, driven by warmer atmospheric temperatures and the intrusion of warm ocean currents in fjords.
+Rising sea level poses a critical threat to coastal communities worldwide, increasing the risk of flooding, shoreline erosion, and habitat loss (Team et al., 2023).
+Estimating its contribution to sea level requires a detailed understanding of subglacial topography, which influences ice flow and basal friction (Barnes et al., 2021; Åkesson et al., 2021).
 However, accurately inferring the bed topography remains a significant challenge due to limited observational coverage, especially in the ice sheet interior.
-Traditional numerical methods, such as mass conservation models and geostatistic techniques (Morlighem et al., 2011; Bamber et al., 2001, 2013), have been instrumental but often struggle with sparse data.
+Traditional mapping methods, such as mass conservation models and geostatistic techniques (Morlighem et al., 2011; Bamber et al., 2001, 2013), have been instrumental but often struggle in regions of slower flow.
 Deep learning (LeCun et al., 2015), with its ability to learn complex relationships from multimodal data, has shown promise in solving geophysical problems.
@@
 The insights gained from this study contribute to a deeper understanding of subglacial dynamics and their implications for sea-level rise projections.
-To summarize, the contributions of this paper are threefold: (i) DeepTopoNet Framework: A novel deep learning framework integrating radar observations and BedMachine data, featuring a CNN architecture enhanced with gradient-based covariates and trend-surface features; (ii) Dynamic Loss-Balancing Mechanism: An adaptive mechanism ensuring robust performance in regions with sparse radar data by balancing contributions from multimodal data sources; and (iii) Comprehensive Evaluation: Extensive benchmarking across Greenland’s Upernavik and unseen regions, demonstrating superior performance over traditional interpolation and machine learning methods.
+To summarize, the contributions of this paper are threefold: (i) DeepTopoNet Framework: A novel deep learning framework integrating radar observations and BedMachine data, featuring a CNN architecture enhanced with gradient-based covariates and trend-surface features; (ii) Dynamic Loss-Balancing Mechanism: An adaptive mechanism ensuring robust performance in regions with sparse radar data by balancing contributions from multimodal data sources; and (iii) Comprehensive Evaluation: Extensive benchmarking across Greenland’s Upernavik Isstrøm sector and unseen regions, demonstrating superior performance over traditional interpolation and machine learning methods.
 2. Related Work
@@
 Subglacial Bed Topography Prediction.
-Inferring the subglacial bed topography has been a critical focus in glaciology due to its implications for ice sheet mass balance and their contributions to global sea-level rise.
+Inferring the subglacial bed topography has been a critical focus in glaciology due to its implications for ice sheet mass balance and its contribution to global sea-level rise.
 Recent advancements have integrated machine learning, physics-based models, and hybrid approaches to tackle this complex problem.
@@
 It employs a mass conservation approach in fast flowing regions, combining radar-derived ice thickness measurements, satellite velocity data, and surface mass balance reconstructions to produce a high-resolution bed topography map.
-The idea of mass conservation is to infer the bed elevation by ensuring conservation of mass: , where is the ice velocity, is surface elevation, is bed elevation, and is the apparent mass balance that combines observed thickness change and surface mass balance from a regional climate model.
+The idea of mass conservation is to infer the bed elevation by ensuring conservation of mass: , where is the apparent mass balance that combines observed thickness change and surface mass balance from a regional climate model.
 BedMachine ensures physically realistic reconstructions in regions of fast-flowing ice, but remains highly uncertain in slower moving regions where it is based on a simple interpolation scheme that connects existing radar data.
@@
 are dynamic weights that adjust during training, ensuring balanced optimization with as a small smoothing factor to avoid division by zero.
+The magnitudes of and are inherently influenced by the proportion of radar-covered () and non-radar () pixels in each mini-batch.
+In radar-dense batches, is computed over many radar points and tends to dominate, resulting in a smaller and a larger .
+In radar-sparse batches, dominates, increasing and reducing .
+This coupling between loss magnitude and data coverage means that the weighting terms implicitly adapt to the local radar density of each batch without requiring explicit density computation or manual tuning.
 The objective is to train a a deep convolutional neural network model with parameters to minimize the total loss , ensuring robust predictions across radar-sparse and radar-dense regions: .
@@
 5.1.1. Dataset
-The dataset used in this study is obtained from the Upernavik Isstrøm region in the Central West of Greenland, encompassing four distinct sub-regions.
+The dataset used in this study is obtained from the Upernavik Isstrøm region in Central West Greenland, encompassing four distinct sub-regions.
 Each sub-region covers a spatial extent of grid cells, representing an area with diverse ice dynamics and varying data density.
@@
 Compared to traditional interpolation methods, such as IDW (Johnston et al., 2001) and RBF (Heryudono and Driscoll, 2010), DeepTopoNet delivers significantly improved performance.
-For instance, while IDW achieves moderate results in some sub-regions (e.g., SSIM of 0.87 in Sub-region II), its RMSE values remain substantially higher (e.g., 101.05 in Sub-region I).
+For instance, while IDW achieves moderate results in some sub-regions (e.g., SSIM of 0.87 in Sub-region II), its RMSE values remain substantially higher (e.g., 101.05 m in Sub-region I).
 Similarly, RBF fails to capture finer spatial details, resulting in higher errors across all metrics.
@@
 Against pixel-based machine learning methods, including RF (Sekulić et al., 2020), MLP-PE (Li et al., 2021) and VAE+XGBoost (Yi et al., 2023), DeepTopoNet exhibits a significant advantage.
-RF struggles to capture complex spatial relationships, as reflected in its relatively high RMSE values (e.g., 99.87 in Sub-region I) and low R2 scores across sub-regions.
+RF struggles to capture complex spatial relationships, as reflected in its relatively high RMSE values (e.g., 99.87 m in Sub-region I) and low R2 scores across sub-regions.
 MLP-PE, despite incorporating positional encoding, performs poorly with high errors and negative R2 in some sub-regions (e.g., Sub-region III), indicating limited generalization.
-In contrast, DeepTopoNet effectively leverages spatial features and multimodal data, achieving an RMSE as low as 19.38 and an SSIM of 0.94 in Sub-region I.
+In contrast, DeepTopoNet effectively leverages spatial features and multimodal data, achieving an RMSE as low as 19.38 m and an SSIM of 0.94 in Sub-region I.
 VAE+XGBoost performs better than MLP-PE and RF but still falls short of DeepTopoNet in all sub-regions.
-While VAE+XGBoost benefits from variational autoencoder-based feature extraction and gradient-boosted decision trees for regression, its RMSE remains significantly higher than DeepTopoNet across all sub-regions (e.g., 108.54 in Sub-region I vs.
-19.38 for DeepTopoNet).
+While VAE+XGBoost benefits from variational autoencoder-based feature extraction and gradient-boosted decision trees for regression, its RMSE remains significantly higher than DeepTopoNet across all sub-regions (e.g., 108.54 m in Sub-region I vs.
+19.38 m for DeepTopoNet).
 When compared to patch-based deep learning models, including U-Net (Ronneberger et al., 2015), U-Net++ (Zhou et al., 2018), U-Net3+ (Huang et al., 2020), and Attention U-Net (Oktay et al., 2018), DeepTopoNet shows a clear improvement.
-Among these baselines, Attention U-Net performs the best, achieving competitive results with an RMSE of 33.33 and SSIM of 0.90 in Sub-region I.
-However, DeepTopoNet consistently surpasses Attention U-Net, achieving a much lower RMSE of 19.38 and a higher SSIM of 0.94 in the same sub-region.
+Among these baselines, Attention U-Net performs the best, achieving competitive results with an RMSE of 33.33 m and SSIM of 0.90 in Sub-region I.
+However, DeepTopoNet consistently surpasses Attention U-Net, achieving a much lower RMSE of 19.38 m and a higher SSIM of 0.94 in the same sub-region.
 This indicates that DeepTopoNet’s combination of enriched input features, dynamic loss balancing, and robust architecture allows it to better capture spatial relationships and generalize effectively across diverse sub-regions.
@@
 Removing these components leads to a substantial decline in performance.
-For instance, in Sub-region IV, RMSE increases drastically to 457.06, and R2 becomes negative (e.g., -57.74), indicating a failure to capture meaningful spatial patterns.
+For instance, in Sub-region IV, RMSE increases drastically to 457.06 m, and R2 becomes negative (e.g., -57.74), indicating a failure to capture meaningful spatial patterns.
 In other sub-regions, metrics such as SSIM and PSNR are lower compared to DeepTopoNet, emphasizing the importance of these features for accurate predictions.
@@
 Retaining gradient and trend surface features while removing stride significantly improves performance compared to the first scenario.
-For example, in Sub-region I, RMSE reduces to 36.26, and R2 increases to 0.96, demonstrating that gradient and trend surface features are vital for capturing spatial variability.
-However, the performance still lags behind DeepTopoNet in certain metrics, such as RMSE in Sub-region III (27.55 vs.
-32.69 for DeepTopoNet).
+For example, in Sub-region I, RMSE reduces to 36.26 m, and R2 increases to 0.96, demonstrating that gradient and trend surface features are vital for capturing spatial variability.
+However, the performance still lags behind DeepTopoNet in certain metrics, such as RMSE in Sub-region III (27.55 m vs.
+32.69 m for DeepTopoNet).
 w/o BedMachine Loss ().
 Removing the BedMachine Loss () significantly degrades the model’s performance, highlighting its critical role in training DeepTopoNet.
-Across all sub-regions, the absence of leads to substantial increases in RMSE, particularly in Sub-region IV, where RMSE jumps to 106.00 and R2 drops to -2.16, indicating a severe loss of predictive accuracy.
+Across all sub-regions, the absence of leads to substantial increases in RMSE, particularly in Sub-region IV, where RMSE jumps to 106.00 m and R2 drops to -2.16, indicating a severe loss of predictive accuracy.
 Additionally, SSIM and PSNR values decline across all sub-regions, reflecting a reduction in structural similarity and topographic detail retention.
@@
 The full DeepTopoNet framework achieves the best results across most metrics and sub-regions.
-For example, in Sub-region I, RMSE is the lowest at 19.38, and SSIM reaches 0.94, indicating excellent structural similarity between predictions and ground truth.
+For example, in Sub-region I, RMSE is the lowest at 19.38 m, and SSIM reaches 0.94, indicating excellent structural similarity between predictions and ground truth.
 This highlights the effectiveness of combining stride, , t, and for robust and accurate bed topography predictions.
@@
 This confirms that is essential for balancing high-resolution radar-derived observations with global-scale topographic trends, ensuring stability and accuracy in bed topography predictions.
-| Prediction/BedMachine (Morlighem et al., 2017, 2020)/Difference |
+| Prediction — BedMachine (Morlighem et al., 2017, 2020) — Difference |
 | IDW (Johnston et al., 2001) |
@@
 As shown in Table 6, DeepTopoNet achieves the best performance across all major evaluation metrics in Sub-region I.
-It produces the lowest MAE (15.90) and RMSE (25.39), while maintaining the highest (0.98), SSIM (0.91), and PSNR (33.90).
+It produces the lowest MAE (15.90 m) and RMSE (25.39 m), while maintaining the highest R2 (0.98), SSIM (0.91), and PSNR (33.90).
 Notably, even though U-Net and U-Net3+ perform moderately well, they fall short in both structural similarity and magnitude-based errors.
@@
 This research has been funded by the NSF HDR Institute for Harnessing Data and Model Revolution in the Polar Regions (iHARP), Award #2118285.
-Code Availability
+Appendix A Code Availability
 The source code for DeepTopoNet and all related experiments is available at: https://github.com/bayuat/DeepTopoNet
+Appendix B Additional Results for Sub-region II, III, and IV
+To further evaluate the performance of the proposed framework and baseline methods, we present additional results for Sub-regions II, III, and IV (see Table 7).
+These regions provide diverse challenges due to variations in radar data density, terrain complexity, and topographic features.
+| Sub-region II — Sub-region III — Sub-region IV |
+| IDW (Johnston et al., 2001) |
+| RBF (Heryudono and Driscoll, 2010) |
+| RF (Sekulić et al., 2020) |
+| MLP-PE (Li et al., 2021) |
+| VAE+XGBoost (Yi et al., 2023) |
+| U-Net (Ronneberger et al., 2015) |